# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

### v1

@@ 2.10. The scattering map
We recall here the scattering map [DdlLS00, DdlLS08].
@@ 2.10.3. Definition of the scattering map
The last hypothesis in Definition  2.23, that ${\Omega_{-}}_{\mid\Gamma}$ is a diffeomorphism from its domain to its range, can always be arranged by restricting $\Gamma$ to a smaller neighborhood where the implicit function theorem applies.
@@
are $\mathcal{C}^{1}$-diffeomorphisms. Thus, we can define the scattering map associated to $\Gamma$ as follows:
@@
In general, the scattering map depends on $\Gamma$ and is only locally defined. In [DdlLS00] there are examples where the local domain of the scattering map cannot be extended to a global one (moving along a cycle in $\Lambda$ leads to lack of monodromy).
The scattering map provides an efficient way to quantify the effect of homoclinic trajectories on the NHIM $\Lambda$. In [DdlLS00, DdlLS06] it is shown that it can be used to study the heteroclinic intersections between invariant objects in $\Lambda$. In [GdlLS20, GLLMS20] it is shown that iterations of the map restricted to the NHIM $\Lambda$ combined with iterates of the scattering map $S$ are closely followed by true orbits.
In typical situations, we have many scattering maps (due to the existence of multiple intersections of the stable and unstable manifolds). All these scattering maps can be used to construct a rich set of orbits.
@@ 3.1. Standing assumptions
Unless otherwise stated, all the results in this section will assume the following:
@@ 3.3. Symplectic properties of scattering maps
In Section 8.2 we give seven different proofs of part (B) of Theorem 3.3. Some of them require slightly different hypotheses. For instance, some of the proofs do not use that $\omega$ is closed or non-degenerate (hence, they apply to non-symplectic contexts), other proofs assume that $\omega$ is $\mathcal{C}^{1}$-bounded, and other ones assume different conditions among the hyperbolic rates and the conformal factor.
@@ 3.4. Results for presymplectic systems
Now we examine the integrability of the kernel of a presymplectic form. For a fixed $x$, the kernel $K_{x}(\omega)$ is a linear subspace of $T_{x}M$, and the family $K_{x}(\omega)$, $x\in M$, determines a distribution on $M$. Rank-11 kernels are just multiples of a vector field and can be integrated by solving ODE’s. For higher rank kernels, the integrability is non-trivial (see [AdlL12]).
@@
That is, the distribution $K$ is closed under taking commutators. This is the hypothesis of Frobenius theorem. A version of Frobenius theorem with low regularity appears in [Har02, pp. 123-124]. See also [Yao23].
@@
Assume that $f$ is conformally presymplectic. If $u\in K_{x}(\omega)$ then, for any $v\in T_{x}M$, we know that $\omega(x)(u,v)=0$ and, consequently:
@@
We can obtain a concrete representation of the dynamics on the space of leaves by taking transversal sections $T_{x},T_{f(x)}$ to the foliation $\mathcal{F}$ at $x$ and $f(x)$, respectively. Each transversal can be endowed with the restriction of $\omega$, which is non-degenerate since the transversal excludes the kernel of the form. Note that $\omega_{\mid T_{x}}$ and $\omega_{\mid T_{f(x)}}$ are closed because the exterior derivative commutes with the restriction. Then, given $y\in T_{x}$, associate to it $\tilde{y}=\tilde{f}(y)\in T_{f(x)}$ defined by $\tilde{y}=H\circ f(y)$ where $H$ is the holonomy map sending $x$ to $f(x)$. If $f$ is conformally presymplectic, then $\tilde{f}$ is conformally symplectic from $(T_{x},\omega_{\mid T_{x}})$ to $(T_{f(x)},\omega_{\mid T_{f(x)}})$. Similar constructions appear in [LM87, p. 106 ff.]. We will apply a similar construction to the scattering map.
@@
The constant rank property of presymplectic forms is taken as part of the definition in some treatments [Sou97, LM87]. As we noted earlier, the constant rank assumption is not an open condition, since adding an arbitrary small perturbation to the form may decrease the dimension of the kernel.
@@ 3.4.1. The scattering map for conformally presymplectic systems
Assume further that the symplectic form is presymplectic exact, i.e., $\omega=d\alpha$. Then, the scattering map $S$ is exact, that is
@@
The proof is given in Section 10.
@@ 4. Results on topology of manifolds with conformally symplectic dynamics
In this section, we show that there are interactions between the (co)homology of the manifold and the set of conformally symplectic factors. In particular, we present an answer to a question posed in [AF24, p. 160].
In this section we assume that we have a well defined cohomology theory and that the $1$ and $2$-cohomology considered are finite dimensional (hence, we can define pull-back operators and they are finite dimensional). Then, we obtain results for maps which are conformally symplectic with respect to forms in this class.
@@ 4.1. Topological obstructions to exactness
Assume that $\omega=d\alpha$ for some $1$-form $\alpha$, and that $\eta$ is not an eigenvalue of $f^{\#1}$ acting on the $1$-cohomology group.
Then, there exists a closed $1$-form $\beta$ so that $\tilde{\alpha}=\alpha+\beta$ satisfies $\omega=d\tilde{\alpha}$ and
@@
for some primitive function $P$, and therefore $f$ is exact with respect to $\tilde{\alpha}$.
The $1$-form $\beta$ above is unique up to the addition of an exact form $dG$ for $G$ a function. The function $P$ is unique up to the addition of $G\circ f-\eta G$ for some globally defined function $G$.
@@
Taking $1$-cohomology on the left-side of (4.1)
we get
(4.2) $f^{*}{\tilde{\alpha}}-\eta{\tilde{\alpha}}]=[f^{*}{\alpha}-\eta{\alpha}]+f^{\#1}[\beta]-\eta[\beta].$
As $\eta$ is not an eigenvalue of $f^{\#1}$, we can find a unique $\beta]$ so that the $1$-cohomology of (4.2) vanishes. This determines $\beta$ up to the addition of the differential of a function $G$. The corresponding change in the primitive function $P$ follows from (2.9). ∎
@@
Some of these theories may lead to trivial obstructions, and several cohomology theories may lead to the same obstruction. In Example 4.14 we will apply simplicial cohomology obtaining interesting obstructions (see Proposition 4.10), and in Remark 4.17 we will apply Borel-Moore homology obtaining only trivial obstructions.
@@
For any map $g$ that is $\mathcal{C}^{1}$-homotopic to $f$, we have that $g^{\#}$ is isomorphic to $f^{\#}$ and hence the conformal factor $\eta$ for a non-exact symplectic form has to remain constant under $\mathcal{C}^{1}$-homotopy.
Under assumption (U1), if $g,f$ are $\mathcal{C}^{1}$-close (hence $\mathcal{C}^{1}$ homotopic) conformally symplectic mappings for the same non-exact form, the conformal factors have to be the same.
Hence for a non-exact form, the set of conformally symplectic factors is locally constant for $f$ in the $\mathcal{C}^{1}$ topology.
@@
A particular case of Remark 4.7 is that, when the manifold has finite dimensional cohomology, symplectic diffeomorphisms for a non-exact form cannot be perturbed into a conformally symplectic. The Example 5.2 presents a different argument for the same phenomenon.
In contrast, for exact forms $\omega=d\alpha$, given a function $H$, we can define vector fields $X$ by $i_{X}\omega=dH+\sigma\alpha$.
It is standard that the time-$t$ map of the flow of $X$ is conformally symplectic for $\omega$ with a conformal factor $\exp(\sigma t)$.
@@
If $f$ is homotopic to the identity, (which is implied by $f$ being the time-$1$ map of a time-dependent flow or by $f$ being $\mathcal{C}^{1}$-close to identity) then the induced maps $f^{\#}$ on $1$ and $2$-cohomology are the identity. If the conformal factor $\eta\neq 1$, then $\eta$ is not an eigenvalue of $f^{\#1}$, $f^{\#2}$ and Lemmas 4.3 and  4.4 apply. In this case we recover [AF24, Proposition 9].
Time-11 maps of conformally symplectic vector fields are exact because there are explicit formulas for the primitive function, [AF24].
@@ 4.2. Topological obstructions to conformal factors
In this section we will address the following question formulated in [AF24, p. 160]:
@@
We will not attempt a general setting as in [AF24], rather we are just presenting some examples in a concrete manifold: the Cartesian product of a torus and an Euclidean space. Since this manifold is a deformation retract of the torus, all the cohomology theories give the same answer. We will also consider diffeomorphisms with bounded derivatives and bounded forms.
@@ 4.2.1. Computation of $f^{\#2}$ on some manifolds
The matrix elements of $f^{\#2}$ can be computed explicitly when $M=\mathbb{R}^{d}\times\mathbb{T}^{d}$, where $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ and $d\geq 2$. We denote the coordinates on $M$ by $(I,\theta)$.
@@
In our case, any conformal factor will be the product of two algebraic numbers of degree $d$. This is related to a question raised in [AF24, p. 165], where a very similar phenomenon is observed in some examples.
@@
Since the action on cohomology remains constant under a homotopy, we obtain that the action of $F^{1}$ on 2-cohomology is the same as that of $F^{0}$. The latter is just $A^{\wedge 2}$ (the wedge product  [^4] We recall that $A^{\wedge 2}$ is defined by $A^{\wedge 2}(\alpha\wedge\beta)=(A\alpha)\wedge(A\beta)$ for all vectors $\alpha,\beta$ of $A$ with itself), which agrees with the direct calculation.
@@ 4.2.2. A concrete example
The main observation is that if $B$ is an automorphism of the torus with simple eigenvalues such that $BA=AB$, then $A$ and $B$ have a common set of eigenvectors555 If $Av=\sigma v$, multiplying by $B$ on the left and commuting, we have $A(Bv)=\sigma(Bv)$, Since we are assuming the space of eigenvectors of $A$ with eigenvalue $\sigma$ is $1$-dimensional, we have that there exists $\nu\in\mathbb{R}$ such that $Bv=\nu v$., even though the eigenvalues may be quite different and, indeed, independent over the rationals. Now, if $A,B$ have the same eigenvectors, then the maps $f_{A},f_{B}$ as in (4.6) are conformally symplectic with factors $\eta_{A},\eta_{B}$ respectively. Therefore, the set $\mathcal{R}$ contains the multiplicative group generated by $\eta_{A},\eta_{B}$. This group could well be dense.
The construction of integer matrices that commute and have simple eigenvalues is not obvious. We just point out that several such examples, with many extra properties, appear in [KN11, p. 61 ff.] motivated by the theory of Abelian actions on the torus. For some of these examples, $\mathcal{R}$ is dense in $\mathbb{R}$.

### v2

@@ 2.10. The scattering map
We recall here the scattering map [DLS00, DLS08].
@@ 2.10.3. Definition of the scattering map
The last hypothesis in Definition  2.22, that ${\Omega_{-}}_{\mid\Gamma}$ is a diffeomorphism from its domain to its range, can always be arranged by restricting $\Gamma$ to a smaller neighborhood where the implicit function theorem applies.
@@
are $\mathcal{C}^{1}$-diffeomorphisms.
We define the scattering map associated to $\Gamma$ as follows:
@@
Since the projections $\Omega_{\pm}$ satisfy (2.24), we have the $\Omega_{\pm}^{\Gamma}$ and $S^{\Gamma}$ satisfy the equivariance relations
(2.31) $\begin{split}\Omega_{\pm}^{\Gamma}=&f_{\mid\Lambda}^{-n}\circ\Omega_{\pm}^{f^{n}(\Gamma)}\circ f^{n},\\ S^{\Gamma}=&f_{\mid\Lambda}^{-n}\circ S^{f^{n}(\Gamma)}\circ f^{n}.\end{split}$
Remark 2.25.
@@
In general, the scattering map depends on $\Gamma$ and is only locally defined. In [DLS00] there are examples where the local domain of the scattering map cannot be extended to a global one (moving along a cycle in $\Lambda$ leads to lack of monodromy).
The scattering map provides an efficient way to quantify the effect of homoclinic trajectories on the NHIM $\Lambda$. In [DLS00, DelshamsLS06a] it is shown that it can be used to study the heteroclinic intersections between invariant objects in $\Lambda$. In [GideaLS20, GideaLS20a] it is shown that iterations of the map restricted to the NHIM $\Lambda$ combined with iterates of the scattering map $S$ are closely followed by true orbits.
In typical situations, we have many scattering maps (due to the existence of multiple intersections of the stable and unstable manifolds). All these scattering maps can be used to generate rich dynamical behavior.
@@ 3. Main results on NHIMs and scattering maps
In this section we state results on NHIMs. The proofs appear in Sections 7 and 8. Other results which do not involve involve NHIMs appear in Section 4.
@@ 3.1. Standing assumptions
Let us summarize the standing assumptions introduced so far:
@@ 3.3. Symplectic properties of scattering maps
In Section 8.2 we give seven different proofs of part (B) of Theorem 3.3. Some of them require slightly different hypotheses. For instance, some of the proofs do not use that $\omega$ is closed or non-degenerate, other proofs assume that $\omega$ is $\mathcal{C}^{1}$-bounded, and other ones assume different conditions among the hyperbolic rates and the conformal factor. Thus, they are also applicable in non-symplectic contexts, as in the applications described in the Appendix D.
@@ 3.4. Results for presymplectic systems
Now we examine the integrability of the kernel of a presymplectic form. For a fixed $x$, the kernel $K_{x}(\omega)$ is a linear subspace of $T_{x}M$, and the family $K_{x}(\omega)$, $x\in M$, determines a distribution on $M$. Rank-11 kernels are just multiples of a vector field and can be integrated by solving ODE’s. For higher rank kernels, the integrability is non-trivial (see [LibermannM87, Souriau97, AlishahL12]).
@@
That is, the distribution $K$ is closed under taking commutators. This is the hypothesis of Frobenius theorem. A version of Frobenius theorem with low regularity appears in [Hartman, pp. 123-124]. See also [Yao23].
@@
Assume that $f$ is conformally presymplectic. If $u\in K_{x}(\omega)$ then, for any $v\in T_{x}M$, $\omega(x)(u,v)=0$ and, consequently:
@@
We can obtain a concrete representation of the space of leaves and the dynamics on it by taking transversal sections $T_{x},T_{f(x)}$ to the foliation $\mathcal{F}$ at $x$ and $f(x)$, respectively. Each transversal can be endowed with the restriction of $\omega$, which is non-degenerate since the transversal excludes the kernel of the form. Note that $\omega_{\mid T_{x}}$ and $\omega_{\mid T_{f(x)}}$ are closed because the exterior derivative commutes with the restriction. Then, given $y\in T_{x}$, associate to it $\tilde{y}=\tilde{f}(y)\in T_{f(x)}$ defined by $\tilde{y}=H\circ f(y)$ where $H$ is the holonomy map sending $x$ to $f(x)$. If $f$ is conformally presymplectic, then $\tilde{f}$ is conformally symplectic from $(T_{x},\omega_{\mid T_{x}})$ to $(T_{f(x)},\omega_{\mid T_{f(x)}})$. Similar constructions appear in [LibermannM87, p. 106 ff.]. We will apply a similar construction to the scattering map.
@@
The constant rank property of presymplectic forms is taken as part of the definition in some treatments [Souriau97, LibermannM87]. As we noted earlier, the constant rank assumption is not an open condition, since adding an arbitrary small perturbation to the form may decrease the dimension of the kernel.
@@ 3.4.1. The scattering map for conformally presymplectic systems
Assume further that the presymplectic form is exact, i.e., $\omega=d\alpha$. Then, the scattering map $S$ is exact, that is
@@
The proof of Theorem 3.9 is given in Section 10.
@@ 4. Results on topology of manifolds with conformally symplectic dynamics
In this section, we show that there are interactions between the (co)homology of the manifold and the set of conformally symplectic factors. In particular, we present an answer to a question posed in [ArnaudF24, p. 160].
We do not need many of the analytical assumptions in Section 3.1, particularly those concerning NHIMs.
We assume that we have a well defined cohomology theory and that the $1$ and $2$-cohomology considered are finite dimensional (hence, we can define pull-back operators and they are finite dimensional). Then, we obtain results for maps which are conformally symplectic with respect to forms in this class.
@@ 4.1. Topological obstructions to exactness
Assume that $\omega=d\alpha$ for some $1$-form $\alpha$.
  • (i)
    If $\eta$ is not an eigenvalue of $f^{\#1}$, then there exists a closed $1$-form $\beta$ so that $\tilde{\alpha}=\alpha+\beta$ satisfies $\omega=d\tilde{\alpha}$ and
@@
    for some primitive function $P$, and therefore $f$ is exact with respect to $\tilde{\alpha}$. The $1$-form $\beta$ is unique up to the addition of an exact form $dG$, for some function $G:M\to\mathbb{R}$. Hence, the function $P$ is unique up to the addition of $G\circ f-\eta G$.
  • (ii)
    If $\eta$ is an eigenvalue of $f^{\#1}$ we have the following dichotomy:
    • (ii.a)
      If $f^{*}\alpha-\eta\alpha]\not\in\textrm{Range}(f^{\#1}-\eta\textrm{Id})$ then there is no closed $1$-form $\beta$ such that $f$ is exact for $\tilde{\alpha}$.
    • (ii.b)
      If $f^{*}\alpha-\eta\alpha]\in\textrm{Range}(f^{\#1}-\eta\textrm{Id})$ then there exists a closed $1$-form $\beta$ such that $f$ is exact for $\tilde{\alpha}$. Moreover, such $\beta$ is unique up to the addition of an exact form $dG$ and of a linear combination of $\beta_{1},\ldots,\beta_{l}$, where $\beta_{i}]$, $i=1,\ldots,l$, are a basis of the space of eigenvectors of $f^{\#1}$ corresponding to $\eta$ and $l$ is the algebraic multiplicity of $\eta$ as an eigenvalue of $f^{\#1}$.
  • (iii)
    If $f$ is symplectic, i.e., $\eta=1$, and $f^{\#1}=\textrm{Id}$, then either $f$ is exact for all actions $\alpha$ or $f$ is not exact for any $\alpha$.
@@
Let $\tilde{\alpha}=\alpha+\beta$, with $\beta$ closed $1$-form.
The map $f$ is exact with respect to $\tilde{\alpha}$ if and only if we have (4.1). Taking $1$-cohomology we get
(4.2) $f^{*}{\tilde{\alpha}}-\eta{\tilde{\alpha}}]=[f^{*}{\alpha}-\eta{\alpha}]+f^{\#1}[\beta]-\eta[\beta]=0,$
which is equivalent to
(4.3) $(f^{\#1}-\eta\textrm{Id})[\beta]=-[f^{*}\alpha-\eta\alpha].$
For part (i), as $\eta$ is not an eigenvalue of $f^{\#1}$, we can find a unique solution $\beta]$ for (4.3). This determines $\beta$ up to the addition of the differential of a function $G$. The corresponding change in the primitive function $P$ follows from (2.9).
Part (ii.a) follows from linear algebra.
For part (ii.b), if $\beta]$ and $\beta^{\prime}]$ are two solutions of (4.3), then $\beta]-[\beta^{\prime}]$ is in the eigenspace of $\eta$.
For part (iii), we have that $\eta=1$ is an eigenvalue of $f^{\#1}=\textrm{Id}$ and $\textrm{Range}(f^{\#1}-\eta\textrm{Id})=\{0\}$.
Then, we observe that for any closed form $\beta$,
$f^{\#1}(\alpha+\beta)-(\alpha+\beta)=[f^{\#1}\alpha-\alpha].$
Therefore, the map $f$ is exact for all $\alpha+\beta$ – all possible action forms – if and only if $f$ is exact for $\alpha$. ∎
Remark 4.5.
The assumption in part (iii) that $f^{\#1}$ is the identity is natural for symplectic maps on compact manifolds. This occurs when one considers a cohomology theory which is invariant under homotopy and $f$ lies in the connected component of the identity in the group of symplectic diffeomorphisms. This is the case, for instance, of the time-$1$ map of an (autonomous) Hamiltonian flow.
The dichotomy in (iii) of $f$ being exact for all actions forms or for none, is not true non-symplectic maps (see Example 5.1), or for symplectic maps not in the connected component of identity.
@@
Some of these theories may lead to trivial obstructions, and several cohomology theories may lead to the same obstruction. In Example 4.15 we will apply simplicial cohomology to obtain obstructions for $\mathcal{R}$. For this example, we show that these obstructions are optimal (see Proposition 4.11). In Remark 4.18 we will apply Borel-Moore homology obtaining only trivial obstructions. The survey [BanyagaHS24] presents some other cohomology theories useful for (locally) conformally symplectic geometry.
@@
Many (co)-homologies theories are homotopy invariant, i.e., homotopic maps induce maps on (co)-homology that are isomorphic [^4] However, this is not the case for the Borel-Moore cohomology; see Remark 4.18. In such a case, the conformal factor $\eta$ for a non-exact symplectic remains constant under $\mathcal{C}^{1}$-homotopy.
Under assumption (U1), if $g,f$ are $\mathcal{C}^{1}$-close (hence $\mathcal{C}^{1}$ homotopic) conformally symplectic mappings for the same non-exact form, then the conformal factors have to be the same.
Hence, for a non-exact form, the set of conformally symplectic factors is locally constant for $f$ in the $\mathcal{C}^{1}$ topology.
@@
A particular case of Remark 4.8 is that a symplectic diffeomorphism for a non-exact form cannot be perturbed into a conformally symplectic one. The Example 5.2 presents a different argument for the same phenomenon.
In contrast, for exact forms $\omega=d\alpha$, given a function $H$, we can define a vector field $X$ by $i_{X}\omega=dH+\sigma\alpha$.
Assuming that the flow of $X$ exists [^5] This is not obvious in unbounded manifolds since $\alpha$ may be unbounded (see Section 9.4) and the standard Cauchy-Picard theorem does not apply. Indeed, there are examples of action forms with solutions that reach infinity in arbitrarily short times., the time-$t$ map of the flow is conformally symplectic for $\omega$ with a conformal factor $\exp(\sigma t)$.
@@
Assume that we consider a cohomology theory that is homotopy invariant. If $f$ is homotopic to the identity, (which is implied by $f$ being the time-$1$ map of a time-dependent flow or by $f$ being $\mathcal{C}^{1}$-close to identity) then the induced maps $f^{\#}$ on $1$ and $2$-cohomology are the identity. If the conformal factor $\eta\neq 1$, then $\eta$ is not an eigenvalue of $f^{\#1}$, $f^{\#2}$ and Lemmas 4.3 and  4.4 (i) apply. In this case, we recover [ArnaudF24, Proposition 9].
Time-11 maps of conformally symplectic vector fields are exact because there are explicit formulas for the primitive function [ArnaudF24].
@@ 4.2. Topological obstructions to conformal factors
In this section we will address the following question formulated in [ArnaudF24, p. 160]:
@@
We will not attempt a general setting as in [ArnaudF24], rather we are just presenting some examples in a concrete manifold: the Cartesian product of a torus and an Euclidean space.
For the obstructions we present, one also needs to specify a cohomology theory. The manifold considered is a deformation retract of the torus, and many homology and cohomology theories agree. But there are others, such as Borel-Moore cohomology (see Remark 4.18) that give different answers. We will use the standard De Rham (co)homology.
Note also that the definition of $\mathcal{R}$ requires specifying a set of allowed diffeomorphisms (e.g., specifying the growth properties at $\infty$). We will consider diffeomorphisms with uniformly bounded derivatives and bounded forms.
@@
We will show, using cohomology, that all elements of $\mathcal{R}$ have to satisfy some conditions. Then, we will provide examples showing that for any number $\eta$ that satisfies the above restrictions, there is a diffeomorphism $f$ that is conformally symplectic for $\omega$ with conformal factor $\eta$.
@@ 4.2.1. Computation of $f^{\#2}$ on some manifolds
The matrix elements of $f^{\#2}$ in the standard De Rham cohomology can be computed explicitly when $M=\mathbb{R}^{d}\times\mathbb{T}^{d}$, where $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ and $d\geq 2$. We denote the coordinates on $M$ by $(I,\theta)$. (Later, in Remark 4.18, we will present results for Borel-Moore theory).
@@
In our case, any conformal factor will be the product of two algebraic numbers of degree $d$. This is related to a question raised in [ArnaudF24, p. 165], where a very similar phenomenon is observed in some examples.
@@
Since the action on de Rham cohomology remains constant under a homotopy, we obtain that the action of $F^{1}$ on 2-cohomology is the same as that of $F^{0}$. The latter is just $A^{\wedge 2}$ (the wedge product  [^6] We recall that $A^{\wedge 2}$ is defined by $A^{\wedge 2}(\alpha\wedge\beta)=(A\alpha)\wedge(A\beta)$ for all vectors $\alpha,\beta$ of $A$ with itself), which agrees with the direct calculation.
@@ 4.2.2. A concrete example
The main observation is that if $B$ is an automorphism of the torus with simple eigenvalues such that $BA=AB$, then $A$ and $B$ have a common set of eigenvectors [^7] If $Av=\sigma v$, multiplying by $B$ on the left and commuting, we have $A(Bv)=\sigma(Bv)$, Since we are assuming the space of eigenvectors of $A$ with eigenvalue $\sigma$ is $1$-dimensional, we have that there exists $\nu\in\mathbb{R}$ such that $Bv=\nu v$., even though the eigenvalues may be quite different and, indeed, independent over the rationals. Now, if $A,B$ have the same eigenvectors, then the maps $f_{A},f_{B}$ as in (4.7) are conformally symplectic with factors $\eta_{A},\eta_{B}$ respectively. Therefore, the set $\mathcal{R}$ contains the multiplicative group generated by $\eta_{A},\eta_{B}$. This group could well be dense.
The construction of integer matrices that commute and have simple eigenvalues is not obvious. We just point out that several such examples, with many extra properties, appear in [KatoKN11, p. 61 ff.] motivated by the theory of Abelian actions on the torus. For some of these examples, $\mathcal{R}$ is dense in $\mathbb{R}$.